20 pts Question 5 Youve lost a lawsuit that requires you to pay $1,500 per month for 120 months starting next month. How mucplease answer
Answer #1

This question requires application of PV of annuity, according to which:

PV =

1- (1+ г)-т Р т Р— Рeriodic Раутent т 3 гate рer period п 3D питber of periods

r = 3%/12 = 0.25% (monthly), n = 120, P = $1500

PV = 1500 * [\frac{1 - (1 + 0.0025)^{-120}}{0.0025}]

PV = 1500 * 103.5618

PV = $155,343